OpenStudy (anonymous):

MEDAL FOR BEST ANSWER! A six-sided number cube with faces numbered 1, 2, 3, 4, 5, and 6 is rolled. What is the probability of rolling a 6 or a number less than 3? A. 1/18 B. 1/6 C. 1/3 D. 1/2

OpenStudy (mathmale):

One (of several) approaches would be to add up the following: P(6) + P(1) + P(2). What are P(6), P(1) and P(2), and why?

OpenStudy (anonymous):

You are trying to find the probability of rolling any 3 specific numbers of a 6 faced cube. 3 is half of 6, therefore the probability of rolling a 1, 2 or 6 is 3/6, which can be simplified to 1/2.

OpenStudy (anonymous):

Very good explanation iaca, thnx!!!

OpenStudy (mathmale):

I basically agree with iaca, but would write out the intermediate steps for additional clarity: P(1) = 1/6 P(2) = 1/6 P(6) = 1/6 P(1 OR 2 OR 6)= 1/6 + 1/6 + 1/6 = 3/6 = 1/2.